OpenStudy (anonymous):

Jahlil is 6 inches shorter than 4 times his sister’s height. Jahlil’s height is 70 inches. Which equation represents h, his sister’s height in inches?

OpenStudy (anonymous):

6 inches SHORTER THAN 4 TIMES his sister's height. Jahlil's height is 70 Shorter than=Less than=(all numbers after it)-6 Times=multiply=4(the value) So the equation would be 4x-6=y x=Jahlil's sister y=Jahlil Jahlil's height is 70 so plug that in for the value that symbolized Jahlil's height and solve. 4x-6=70 +6 +6 4x=76 /4 /4 x=19